# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

### v2

@@ 1 Introduction
Moreover, to ensure scalability with respect to the number of classes, DitHub shares a subset of parameters across different adaptation modules, enhancing memory efficiency without compromising individual class adaptability. At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and Open-Vocabulary capabilities. A high-level overview of our approach is depicted in Footnote 2.
@@ 3 Preliminaries
$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
where $\ell$ is the Object Detection loss function. The term $(\Theta,\Delta)$ represents the set of adapted model parameters. Additionally, $\ell_{\text{zero-shot}}$ quantifies the performance difference in the zero-shot setting before and after adaptation.
@@
LoRA. Given a layer with weight $W\in\mathbb{R}^{d\times k}$, Low-Rank Adaptation [13] (LoRA) uses low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ – with $r\ll\min(d,k)$ – such that the adapted weight $W^{\prime}=W+\Delta W=W+BA$. The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to full fine-tuning.
@@ 4 DitHub
Version Control Systems are ubiquitous tools used to manage and track changes in code repositories. Inspired by this, we propose DitHub — a framework that operates akin to a Version Control System, enabling the creation and expansion of a library of efficient adaptation modules for Vision-Language object detectors. DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Sec. 4.1).
Following [4], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection and use LoRA for fine-tuning. We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Sec. 4.2). Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec. 4.3. The pseudocode of DitHub is provided in Alg. 1.
@@ 4.1 Training Class-Specific Modular Detectors
As better highlighted in Alg. 1, we overcome this issue by introducing an earlier warmup phase at the beginning of each task. This phase is meant to provide a robust, class-agnostic initialization, ensuring that the subsequent specialization stage is more effective.
Warmup. At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig. 2, yellow arrow) to capture class-agnostic yet task-specific knowledge. This phase, run for a fixed number of epochs, provides a common initialization for the $A$ modules before specialization. By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [32, 38]. Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Sec. 4.3. Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
Specialization. After the warmup phase, we proceed to the core of DitHub: the specialization phase. To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig. 2), each dedicated to a single class. These branches undergo independent optimization, represented by a blue arrow in Fig. 2, to capture class-specific capabilities. In doing so, for each training image, we employ the stochastic strategy introduced in Sec. 4.1 to randomly select a single expert for training.
@@
              $A_{c}^{\text{cur}}\leftarrow$ use Eq. 2 $\triangleright$  Merge
@@
To favor selective updates, DitHub performs a fetch operation (Fig. 2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Fig. 2) . For instance, considering the class “dog”, the previously stored expert module $A_{\text{dog}}^{\text{old}}$ is fetched and merged with the current task’s warmup matrix $A_{\text{wu}}$ following a weighted strategy:
@@
Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase. Details on the impact of $\lambda_{A}$ are in Sec. 5.2. Crucially, our merging process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information, resulting in $A_{\text{dog}}^{\text{cur}}$, which serves as the starting point for the specialization phase.
@@ 4.2 Bridging Modularity with Memory Efficiency
To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class. As shown in Fig. 3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows. Sharing the $B$ matrix reduces this requirement by half  [^3] In Sec. A.2, we present DitHub’s results when $B$ is not shared.. Additionally, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder alone (line [Uncaptioned image]). This makes DitHub’s memory demands manageable for practical applications. Moreover, memory efficiency is further optimized by fine-tuning a targeted portion of the base architecture (Sec. A.1).
@@
Unlike Eq. 2, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state. This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Sec. 5.2.
@@ 5.1 Incremental Vision-Language Object Detection
One issue of ODinW-13 is that most classes are confined to a single task among the $13$. However, DitHub is conceived to handle recurring classes across domains, necessitating updates to their respective modules. To better assess incremental detection and module reuse, we introduce ODinW-O (Overlapped), a subset of ODinW-35 that is designed to reflect realistic scenarios where classes can reoccur across tasks over time. Further details are provided in App. C.
@@
Implementation details. We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec. A.1. The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$. The choice of $\lambda$ and its underlying rationale are further discussed in Sec. 5.2. During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see App. E for further hyperparameter details.
Performance on ODinW-13. Tab. 1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO). We evaluate performance in both the full-data setting and few-shot scenarios with $1$, $5$, and $10$ shots per class. All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [4]. Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec. D.1.
@@
Performance on ODinW-O. The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios. As detailed in Sec. C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks. In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
@@
The results in Tab. 2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa. We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts. This design prevents unintentional overwriting of knowledge across classes, tasks, and domains.
Performance analysis as tasks progress. In Fig. 4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task. The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in App. B). On ZCOCO, DitHub exhibits stable performance across tasks, with a negligible loss in zero-shot capabilities — unlike ZiRa, which shows a more pronounced decline. When evaluating performance across the sequence of downstream tasks, both methods demonstrate a decreasing trend, indicating that forgetting remains a significant challenge in Incremental Object Detection. Nevertheless, as more tasks are introduced, the performance gap widens, with DitHub consistently outperforming its competitor. This highlights the effectiveness of our modular design, which enables scalable knowledge organization and adaptation as the number of tasks grows. Moreover, the gap is still present in the ODinW-O setting introduced in this work, which models the realistic scenario where subsequent tasks may share classes. This setting, being inherently more susceptible to forgetting, further underscores the advantages of our approach.
@@ 5.2 Model Analysis and Ablation Studies
We examine the properties of our specialization framework and the benefits of composing class-specific experts. To this aim, we introduce an ablative approach, Experts-not-Experts (EnE), which removes class specialization. EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Sec. 4.1. Instead, given an image from the current task, EnE performs the training step on one module chosen at random among all task classes. As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class. Consequently, they become generalist learners rather than class-specific experts.
In Tab. 3, we compare DitHub with its non-specialized variant, i.e., EnE. Notice that, at inference, EnE acts as an ensemble of class-agnostic modules. In contrast, DitHub employs class-specialized counterparts with an expertise on data containing the same class. This targeted specialization yields a $\boldsymbol{\texttt{+}1.23}$ mAP improvement on average, with additional gains also on the ZCOCO metric. These results affirm that a system of modular experts, each with distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
@@
Figure 5: Results on rare classes of ODinW-13.
Figure 6: Showing unlearning skills of DitHub.
Evaluation on rare classes. As discussed in Sec. 4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their specific use case — including the option to use a single module specialized for a target class. To assess this, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module. Results in Fig. 6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
Specialized modules allows training-free unlearning. DitHub’s advantages go beyond additive Incremental Learning. Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model. This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 33]; to the best of our knowledge, we are the first to explore it within the context of Object Detection. For example, this enables the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, whit $\alpha=0.3$. Fig. 6 illustrates this qualitatively: on the left, the zero-shot model correctly detects common classes, while on the right, subtracting a class module reduces its mAP, effectively unlearning it and leading to false negatives. A quantitative analysis is provided in Sec. D.2.
Sensitivity Analysis of $\boldsymbol{\lambda}$. Varying the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our method. To this end, we define an aggregate metric as the harmonic mean of DitHub’s full-shot performance, balancing specialization (Avg) and zero-shot generalization (ZCOCO). Fig. 8 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Eq. 2). From Fig. 8, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective. Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases. The contrast reflects dataset structure: ODinW-O includes overlapping classes, emphasizing knowledge retention, whereas ODinW-13 has varied class distributions, demanding stronger integration of domain-specific warmup knowledge.
For the incremental adaptation of the shared $B$ matrix, a different trend emerges. The harmonic mean increases with $\lambda_{B}$, peaking at $0.7$ for both datasets. This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Eq. 3, is crucial. Our intuition, supported by [29], is that the latest $B^{\text{opt}}$ matrix integrates knowledge from both earlier and recent tasks, making it more valuable than $B_{t-1}$, which is trained solely on earlier tasks.
@@
Figure 8: Ablative study on the DitHub components. Results are expressed as average performance across the tasks of ODinW-13.
@@
On Warmup and Shared $\boldsymbol{B}$. To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Fig. 8. We establish the zero-shot performance of Grounding DINO (GD) as a reference. Our baseline fine-tuning procedure (Base) trains class-specific $A$ matrices and the shared $B$ matrix independently for each new task, omitting both parameter merging and the warmup phase. While this baseline suffers from complete catastrophic forgetting, it still demonstrates improved performance compared to the zero-shot Grounding DINO.
@@
Building on this baseline, we first add the warmup phase (+wu), which yields a notable improvement by stabilizing training and facilitating the learning of rare classes through robust initialization. Subsequently, incorporating Eq. 3 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters. Crucially, the most substantial performance gain is achieved upon integrating Eq. 2 (DitHub in Fig. 8), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning. Collectively, the synergistic combination of these three key components – the warmup phase for robust initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix for efficiency – yields a substantial $\boldsymbol{\texttt{+}6.13}$ mAP improvement compared to the baseline model.

### v3

@@ 1 Introduction
Moreover, to ensure scalability with respect to the number of classes, DitHub shares a subset of parameters across different adaptation modules, enhancing memory efficiency without compromising individual class adaptability. At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and Open-Vocabulary capabilities. A high-level overview of our approach is depicted in Footnote˜2.
@@ 3 Preliminaries
$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;\Theta+\Delta),\,y)+\ell_{\text{zero-shot}},$ (1)
where $\ell$ is the Object Detection loss function. The term $\Theta+\Delta$ represents the set of adapted model parameters. Additionally, $\ell_{\text{zero-shot}}$ quantifies the performance difference in the zero-shot setting before and after adaptation.
@@
LoRA. Low-Rank Adaptation [13] is a technique to significantly reduce the number of trainable parameters. Given a generic layer with weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ — with $r\ll\min(d,k)$ — such that the adapted weight $W^{\prime}$ is:
$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=BA.$ (2)
The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to fine-tuning the entire model.
@@ 4 DitHub
Version Control Systems are ubiquitous tools used to manage and track changes in code repositories. Inspired by this, we propose DitHub — a framework that operates akin to a Version Control System, enabling the creation and expansion of a library of efficient adaptation modules for Vision-Language object detectors. DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Section˜4.1).
Following [4], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection and use LoRA for fine-tuning. We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Section˜4.2). Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Section˜4.3. The pseudocode of DitHub is provided in Algorithm˜1.
@@ 4.1 Training Class-Specific Modular Detectors
As better highlighted in Algorithm˜1, we overcome this issue by introducing an earlier warmup phase at the beginning of each task. This phase is meant to provide a robust, class-agnostic initialization, ensuring that the subsequent specialization stage is more effective.
Warmup. At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Figure˜2, yellow arrow) to capture class-agnostic yet task-specific knowledge. This phase, which runs for a fixed number of epochs, provides a common initialization for the $A$ modules before their specialization. By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [33, 39]. Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Section˜4.3. Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
Specialization. After the warmup phase, we proceed to the core of DitHub: the specialization phase. To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Figure˜2), each dedicated to a single class. These branches undergo independent optimization, represented by a blue arrow in Figure˜2, to capture class-specific capabilities. In doing so, for each training image, we employ the stochastic strategy introduced in Section˜4.1 to select a single expert for training.
@@
     $A_{c}^{\text{cur}}\leftarrow$ use Equation˜3 $\triangleright$  Merge
@@
To favor selective updates, DitHub performs a fetch operation (Figure˜2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Figure˜2) . For instance, considering the class “dog”, the previously stored expert module $A_{\text{dog}}^{\text{old}}$ is fetched and merged with the current task’s warmup matrix $A_{\text{wu}}$ following a weighted strategy:
@@
Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase. Details on the impact of $\lambda_{A}$ are in Section˜5.2. Crucially, our merging process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information, resulting in $A_{\text{dog}}^{\text{cur}}$, which serves as the starting point for the specialization phase.
@@ 4.2 Bridging Modularity with Memory Efficiency
To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class. As shown in Figure˜3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows. Sharing the $B$ matrix reduces this requirement by half  [^3] In Section A.2, we present DitHub’s results when $B$ is not shared.. Additionally, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder alone (line [Uncaptioned image]). This makes DitHub’s memory demands manageable for practical applications. Moreover, memory efficiency is optimized by fine-tuning a targeted portion of the base architecture (Section˜A.1).
@@
Unlike Equation˜3, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state. This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Section˜5.2.
@@ 5.1 Incremental Vision-Language Object Detection
One issue of ODinW-13 is that most classes are confined to a single task among the $13$. However, DitHub is conceived to handle recurring classes across domains, necessitating updates to their respective modules. To better assess incremental detection and module reuse, we introduce ODinW-O (Overlapped), a subset of ODinW-35 that is designed to reflect realistic scenarios where classes can reoccur across tasks over time. Further details are provided in Appendix˜C.
@@
Implementation details. We focus on Grounding DINO and adapt only its encoder network, as further validated in Section˜A.1. The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$. The choice of $\lambda$ and its underlying rationale are further discussed in Section˜5.2. During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see Appendix˜E for further details.
Performance on ODinW-13. Table˜1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO). We evaluate performance in both the full-data setting and few-shot scenarios with $1$, 55, and $10$ shots per class. All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [4]. Each number represents the average of three runs with different random seeds, with standard deviations provided in Section˜D.1.
@@
Performance on ODinW-O. The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios. As detailed in Section˜C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks. In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
@@
The results in Table˜2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa. We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts. This design prevents unintentional overwriting of knowledge across classes, tasks, and domains.
Performance analysis as tasks progress. In Figure˜4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task. The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in Appendix˜B). On ZCOCO, DitHub exhibits stable performance across tasks, with a negligible loss in zero-shot capabilities — unlike ZiRa, which shows a more pronounced decline. When evaluating performance across the sequence of downstream tasks, both methods demonstrate a decreasing trend, indicating that forgetting remains a significant challenge in Incremental Object Detection. Nevertheless, as more tasks are introduced, the performance gap widens, with DitHub consistently outperforming its competitor. This highlights the effectiveness of our modular design, which enables scalable knowledge organization and adaptation as the number of tasks grows. Moreover, the gap is still present in the ODinW-O setting introduced in this work, which models the realistic scenario where subsequent tasks may share classes. This setting, being inherently more susceptible to forgetting, further underscores the advantages of our approach.
@@ 5.2 Model Analysis and Ablation Studies
We examine the properties of our specialization framework and the benefits of composing class-specific experts. To this aim, we introduce an ablative approach, Experts-not-Experts (EnE), which removes class specialization. EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Section˜4.1. Instead, given an image from the current task, EnE performs the training step on one module chosen at random among all task classes. As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class. Consequently, they become generalist learners rather than class-specific experts.
In Table˜3, we compare DitHub with its non-specialized variant, i.e., EnE. Notice that, at inference, EnE acts as an ensemble of class-agnostic modules. In contrast, DitHub employs class-specialized counterparts with an expertise on data containing the same class. This targeted specialization yields a $\boldsymbol{\texttt{+}1.23}$ mAP improvement on average, with additional gains also on the ZCOCO metric. These results affirm that a system of modular experts, each with distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
@@
Figure 5: Performance comparison of Grounding DINO, EnE, and DitHub on rare object categories from the ODinW-13 dataset.
@@
Evaluation on rare classes. As discussed in Section˜4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their use case — including the option to use a single module specialized for a target class. To assess this, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module. Results in Figure˜6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
Sensitivity Analysis of $\boldsymbol{\lambda}$. Varying the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our method. To this end, we define an aggregate metric as the harmonic mean of DitHub’s full-shot performance, balancing specialization (Avg) and zero-shot generalization (ZCOCO). Figure˜6 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Equation˜3). From Figure˜6, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective. Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases. The contrast reflects dataset structure: ODinW-O includes overlapping classes, emphasizing knowledge retention, whereas ODinW-13 has varied class distributions, demanding stronger integration of domain-specific warmup knowledge.
For the incremental adaptation of the shared $B$ matrix, a different trend emerges. The harmonic mean increases with $\lambda_{B}$, peaking at $0.7$ for both datasets. This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Equation˜4, is crucial. Our intuition, supported by [30], is that the latest $B^{\text{opt}}$ matrix integrates knowledge from both earlier and recent tasks, making it more valuable than $B_{t-1}$, which is trained solely on earlier tasks.
Method Memory (MB) Avg (mAP)
DitHub $r=1$ $\sim 9$ $57.04$
ZiRa $\sim 18$ $57.98$
DitHub $r=2$ $\sim 18$ $60.26$
DitHub $r=4$ $\sim 37$ $60.53$
DitHub $r=8$ $\sim 74$ $61.93$
DitHub $r=16$ $\sim 147$ $62.19$
G-Dino (encoder only) $\sim 659$ $46.80$
Figure 7: (left) Ablation study on DitHub components, showing average performance on ODinW-13 alongside the ZCOCO metric. (right) Memory usage (MB) versus performance (mAP) for DitHub at varying ranks, compared against ZiRa, with performance averaged over the ODinW-13 tasks.
@@
On Warmup and Shared $\boldsymbol{B}$. To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Figure˜7 (left). We establish the zero-shot performance of Grounding DINO (GD) as a reference. Our baseline fine-tuning procedure (Base) trains class-specific $A$ matrices and the shared $B$ matrix independently for each new task, omitting both parameter merging and the warmup phase. While this baseline suffers from complete catastrophic forgetting, it still demonstrates improved performance compared to the zero-shot Grounding DINO.
@@
Building on this baseline, we first add the warmup phase (+wu), which yields a notable improvement by stabilizing training and facilitating the learning of rare classes through robust initialization. Subsequently, incorporating Equation˜4 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters. Crucially, the most substantial performance gain is achieved upon integrating Equation˜3 (DitHub in Figure˜7 (left)), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning. Collectively, the synergistic combination of these three key components – the warmup phase for robust initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix for efficiency – yields a substantial $\boldsymbol{\texttt{+}6.13}$ mAP improvement compared to the baseline model.
Finally, we note that results show a progressive decrease in the ZCOCO metric as the average mAP rises — an expected trade-off as specialization degrades zero-shot capabilities. Notably, the Base fine-tuning baseline scores higher on ZCOCO than the original Grounding DINO zero-shot model. We attribute this counter-intuitive result to catastrophic forgetting. Our full method, DitHub, is trained to be robust across the varied domains of ODinW. In contrast, the Base procedure completely forgets such broad knowledge and overfits to the new tasks. This process of forgetting diverse, out-of-distribution data results in a model that is coincidentally more aligned with the data distribution of the COCO benchmark, leading to an artificially inflated ZCOCO score.
Analysis of the Performance-Memory Tradeoff. To study the trade-off between performance and memory, we examined how the average mAP on ODinW-13 is affected by the LoRA rank $r$, i.e., the dimensionality of the low-rank update matrices. The results, presented in Figure˜7 (right), demonstrate a graceful degradation in performance even as the rank is substantially reduced. Notably, a rank of $r=2$ achieves performance within $2$ mAP points of our best model ($r=16$) while reducing the per-class parameters and memory usage by a factor of eight. We also evaluated the extreme case of $r=1$, which collapses the adaptation matrix into a vector, and found that the performance remains remarkably strong. We attribute this robustness to our framework’s granularity; since each module only encodes the knowledge of a single class, it can operate effectively with fewer parameters than approaches that aggregate information from multiple classes into one module.
In terms of memory efficiency, our method with $r=2$ requires approximately the same memory as ZiRa while performing significantly better ($\boldsymbol{\texttt{+}2.28}$ mAP). Furthermore, our $r=1$ variant delivers highly competitive results, scoring less than $1$ mAP point below ZiRa while using only half the memory. A detailed breakdown of the per-task mAP for these varying ranks is provided in Section˜A.3.
Specialized modules allow training-free unlearning. The advantages of DitHub go beyond additive Incremental Learning. Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model. This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 34]; to the best of our knowledge, we are the first to explore its application within the context of Object Detection. Such a mechanism allows the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, where we set $\alpha=0.3$. Figure˜8 qualitatively demonstrates DitHub’s class-unlearning capability. Each column illustrates the effect of removing a specific class module. For instance, when the horse module is subtracted, the previously detected horses are no longer recognized, resulting in false negatives that indicate successful forgetting. This highlights DitHub’s ability to selectively unlearn target classes while preserving knowledge of unrelated ones. A quantitative analysis is provided in Section˜D.2.
Figure 8: Qualitative results of DitHub’s class unlearning. Each column compares zero-shot Grounding DINO predictions (left) with the output after subtracting a specific class module (right).